OpenStudy (anonymous):

an = [(ln(n))^7]/sqrt(n). Find the limit if it converges, indicate divergence if it diverges. How do I solve this?

OpenStudy (anonymous):

Consider the end behavior of the continuous function\[f(x)=\frac{(\ln(x))^7}{\sqrt{x}}\]Numerator and denominator are both divergent to positive infinity, so use L'Hopital's rule.

OpenStudy (anonymous):

After applying L'Hopital's rule, I got \[(7(\ln x)^62\sqrt{x})/x\] Am I deriving incorrectly?

OpenStudy (anonymous):

Pretty close to what I got, but I didn't simplify the denominator. If you move the 1/x term to the denominator, your answer looks like\[f_1(x)=\frac{7(\ln(x))^6}{\sqrt{x}}\]Now repeat L'Hopital's rule a few more times, and you will eventually get a constant divided by the square root of x. I think that probably converges, as does your sequence.

OpenStudy (experimentx):

after you apply L'hospital rule 7 times you will get \( \large \frac{7*6*5*4*3*2*1}{(1/2)^7sqrt(x)}\)

OpenStudy (anonymous):

OOOPS!! Dropped a factor in the derivative of the denominator. Hate it when that happens. Regardless, it doesn't change the end behavior of the continuous function, or the convergence of the sequence.

OpenStudy (anonymous):

aqua1995, you've got this, right?

OpenStudy (anonymous):

I'm still confused... I thought you had to take the derivative of the denominator was well when applying L'Hospital's Rule. Wouldn't that give you (1/2)n^(-1/2) in the denominator?

OpenStudy (anonymous):

Yeah, it would, but there is a factor of 1/x in the numerator. If we move that term to the denominator, you get sqrt(x) there. We can move the factor of 1/2 to the numerator as two, also (that is the factor that I goofed on earlier). So after seven repetitions of L'Hopital's rule, we get \[f_7(x)=\frac{2^7*7!}{\sqrt{x}}\]which converges.

OpenStudy (anonymous):

See it now?

OpenStudy (anonymous):

I see where you're going with the 2, but would the denominator be 1/sqrt(x) and not just sqrt(x)?

OpenStudy (anonymous):

No. When we take 1/x in the numerator, the one (factor) stays in the numerator, and the x moves to the denominator. With the existing factor of 1/sqrt(x), it becomes sqrt(x).

OpenStudy (anonymous):

Hope I was helpful. I've gotta go to class now.

OpenStudy (anonymous):

Ahh, I think I got it. Thanks so much! :-)

OpenStudy (anonymous):

So this means it converges to zero?

OpenStudy (anonymous):

Yup.